### v1

# PHENIX Collaboration

Abstract
The PHENIX experiment has performed a systematic study of identified charged-hadron ($\pi^{\pm}$, $K^{\pm}$, $p$, $\bar{p}$) production at midrapidity in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}=193$ GeV. Identified charged-hadron invariant transverse-momentum ($p_{T}$) and transverse-mass ($m_{T}$) spectra are presented and interpreted in terms of radially expanding thermalized systems. The particle ratios of $K/\pi$ and $p/\pi$ have been measured in different centrality ranges of large (Cu$+$Au, U$+$U) and small ($p$$+$Al, ${}^{3}$e$+$Au) collision systems. The values of $K/\pi$ ratios measured in all considered collision systems were found to be consistent with those measured in $p$$+$$p$ collisions. However the values of $p/\pi$ ratios measured in large collision systems reach the values of $\approx 0.6$, which is $\approx 2$ times larger than in $p$$+$$p$ collisions. These results can be qualitatively understood in terms of the baryon enhancement expected from hadronization by recombination. Identified charged-hadron nuclear-modification factors ($R_{AB}$) are also presented. Enhancement of proton $R_{AB}$ values over meson $R_{AB}$ values was observed in central ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The proton $R_{AB}$ values measured in $p$$+$Al collision system were found to be consistent with $R_{AB}$ values of $\phi$, $\pi^{\pm}$, $K^{\pm}$, and $\pi^{0}$ mesons, which may indicate that the size of the system produced in $p$$+$Al collisions is too small for recombination to cause a noticeable increase in proton production.
I INTRODUCTION
The quark-gluon plasma (QGP) [1, 2] is a state of matter that exists at extremely high temperature and density. The QGP comprises deconfined, strongly interacting quarks and gluons, which are ordinarily confined inside hadrons. Quantum chromodynamics, the theory of the strong nuclear force, predicts QGP formation in high-energy collisions of heavy nuclei. The QGP existence have been verified by experimental observations of QGP signatures [1, 3], such as strangeness enhancement [4, 5, 6, 7, 8, 9], jet quenching [10, 11, 12, 13], and baryon enhancement [14, 15, 16, 17, 18, 19].
The QGP is formed in heavy ion collisions (such as Cu$+$Au, U$+$U) at temperatures larger than 300 MeV [20, 21], but expands and cools down in a short time ($\approx$10 fm/c [3]). When the critical temperature ($\approx 175$ MeV [1, 20]) is reached, the hadronization process begins. It was expected [20] that in relativistic ion collisions protons are produced roughly 3 times less than pions, reflecting the mass difference between them and the requirement for a nonzero baryon number to form a proton. Indeed, the values of the proton to pion ratio ($p/\pi$) measured by the PHENIX experiment in $p$$+$$p$ collisions do not exceed a value of 0.3. However, in central Au$+$Au collisions, the $p/\pi$ ratio reaches a value of 0.8, indicating that protons and $\pi$-mesons are produced in nearly equal proportion. The anomalous proton production observed by the PHENIX experiment [22] in Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ =130 GeV [15] and later at $\sqrt{s_{{}_{NN}}}$ =200 GeV [16] was successfully described by the recombination model. The recombination model [23, 24] is a hadronization model that considers hadron formation as a result of combining quarks that are nearby in phase space.
Small collision systems (such as $p$$+$Al, ${}^{3}$e$+$Au) were previously studied primarily to investigate cold-nuclear-matter effects [25, 26] (Cronin enhancement [27, 28], multiple-parton scattering [29], modifications of the initial nuclear-parton distribution functions [30, 31]). It was thought [32, 25] that the size and lifetime of the system produced in small collision systems are not sufficient for QGP formation. However, in 2019 the PHENIX experiment reported on the observation of elliptic and triangular flow patterns of charged particles produced in $p$/$d$/${}^{3}$e$+$Au collisions [32]. The results of these measurements have been described in the frame of hydrodynamical models, which include the formation of short-lived QGP droplets. Comparison of identified charged-hadron production in small and large collision systems may enable systematic studies to determine the minimal conditions of QGP formation and to investigate influence of collision geometry and system size on hadron production.
This paper presents new data on identified charged-hadron production in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Discussed are the influence of collision centrality and geometry on identified charged-hadron invariant $p_{T}$ and $m_{T}$ spectra, nuclear-modification factors, and ratios.
II DATA ANALYSIS
The data sets used in the present analysis were collected by the PHENIX experiment during the operational periods of the Relativistic Heavy Ion Collider in calendar years 2012 (Cu$+$Au and U$+$U), 2014 (${}^{3}$e$+$Au), and 2015 ($p$$+$Al). The PHENIX experiment [22] comprised several subdetectors that are grouped into four main blocks called arms. The central arms—east and west—cover midrapidity ($|\eta|<0.35$) and are intended for measurement of electrons, photons, and charged hadrons. In the present analysis, the drift chambers (DC) [33] are used for momentum determination, the time-of-flight (TOF) wall [34] is used for charged hadron identification, and the beam-beam counters (BBC) [35] are used for centrality determination. Detailed information about the PHENIX experiment can be found in Refs. [22, 36, 35].
II.1 Event Selection
The minimum-bias (MB) trigger was used for event selection. The MB trigger selects events that correspond to simultaneous signals in the north (forward rapidity) and south (backward rapidity) BBC detectors. The collision vertex coordinate is required to be $|z_{{\rm vtx}}|<30~{}{\rm cm}$ relative to the center of the spectrometer. Determination of centrality in the PHENIX experiment is done with the BBC according to the procedure described in [37]. The numbers of participating nucleons $\langle N_{\rm part}\rangle$ as well as numbers of binary nucleon-nucleon collisions $\langle N_{\rm coll}\rangle$ are calculated using Glauber Monte Carlo simulations. The $\langle N_{\rm part}\rangle$ and $\langle N_{\rm coll}\rangle$ values are presented in the Table 2.
Table 1: Summary of the $\langle N_{\rm coll}\rangle$ and $\langle N_{\rm part}\rangle$ values calculated using Glauber Monte Carlo simulation.
Table 2: The $p_{T}$ ranges (GeV/$c$) of identified charged-hadron yields measurements.
System Centrality $\langle N_{\rm coll}\rangle$ $\langle N_{\rm part}\rangle$
$p$$+$Al 0%–72% 2.1$\pm$0.2 3.1$\pm$0.1
0%–20% 3.4$\pm$0.3 4.4$\pm$0.3
20%–40% 2.3$\pm$0.2 3.3$\pm$0.1
40%–72% 1.7$\pm$0.1 1.6$\pm$0.2
${}^{3}$e$+$Au 0%–88% 10.4$\pm$0.7 11.3$\pm$0.5
0%–20% 22.3$\pm$1.7 21.1$\pm$1.3
20%–40% 14.8$\pm$1.1 15.4$\pm$0.9
40%–60% 8.4$\pm$0.6 9.5$\pm$0.6
0%–88% 3.4$\pm$0.3 4.8$\pm$0.3
Cu$+$Au 0%–80% 123.8$\pm$12.0 70.4$\pm$3.0
0%–20% 313.8$\pm$28.4 154.8$\pm$4.1
20%–40% 129.3$\pm$12.4 80.4$\pm$3.3
40%–60% 41.8$\pm$5.3 34.9$\pm$2.9
60%–80% 10.1$\pm$2.0 11.5$\pm$1.8
U$+$U 0%–80% 342$\pm$30 143$\pm$5
0%–20% 935$\pm$98 330$\pm$6
20%–40% 335$\pm$33 259$\pm$7
40%–60% 81$\pm$13 65$\pm$6
60%–80% 17$\pm$4 18$\pm$3
Hadron $p$$+$Al ${}^{3}$e$+$Au Cu$+$Au U$+$U
$\pi^{+}$ 0.5–2.0 0.5–3.0 0.5–3.0 0.5–3.5
$\pi^{-}$ 0.5–2.0 0.5–2.0 0.5–3.0 0.5–3.0
$K^{+}$ 0.5–1.8 0.5–2.0 0.5–2.0 0.5–2.0
$K^{-}$ 0.5–1.8 0.5–2.0 0.5–2.0 0.5–2.0
$p$ 0.5–4.0 0.5–4.0 0.5–4.0
$\bar{p}$ 0.5–2.5 0.5–4.0 0.5–4.0 0.5–4.0
Table 2: The $p_{T}$ ranges (GeV/$c$) of identified charged-hadron yields measurements.
II.2 Particle Identification
Charged hadrons are detected in the TOF wall and the DC of the PHENIX experiment. The squared mass of tracks can be determined in accordance with Eq. 1:
$m^{2}=\frac{p^{2}}{c^{2}}\left(\frac{t^{2}c^{2}}{L^{2}}-1\right),$ (1)
where $p$ is the momentum of the particle measured with DC, $L$ is the particle flight path-length from the event vertex to the TOF detector, $t$ is the time-of-flight, measured in TOF detector, $c$ is the speed of light [36]. Table 2 presents $p_{T}$ ranges of identified charged-hadron-yield measurements.
The distribution of $m^{2}$ multiplied by charge versus transverse momentum ($p_{T}$) as found from TOF timing, DC momentum, and the path length is presented in Fig.1. Signals corresponding to $\pi^{\pm}$, $K^{\pm}$, and $p$ and $\bar{p}$ are clearly distinguishable. In $p$$+$Al collisions, the proton sample has a large contamination of spallation protons from the inner silicon detectors. For that reason, they are excluded from measurements that require absolute normalization, like the spectra, ratios, and nuclear modification factors reported in this manuscript.
[…]
$m_{\rm TOF}^{2}(p_{T}),\sigma_{\rm TOF}(p_{T})=p_{0}+\frac{p_{1}}{p_{T}}+\frac{p_{2}}{p_{T}^{2}}+p_{3}\cdot e^{\sqrt{p_{T}}}+p_{4}\cdot\sqrt{p_{T}},$ (2)
where $p_{0}$, $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are fit parameters, and the set of fit parameters is different for $m_{\rm TOF}^{2}(p_{T})$ and $\sigma_{\rm TOF}(p_{T})$.
Figure 1: Distribution of hadron squared mass multiplied by charge ($m^{2}\cdot{\rm charge}$) vs. hadron $p_{T}$ as found from TOF timing, DC momentum, and the path length. Black solid lines represent PID cuts, based on Eq. 2, which were used for hadron identification.
II.3 Corrections to the raw data
The measured values of identified charged hadron primary yields should be corrected for the geometric acceptance of the detectors, detector efficiency, and analysis cuts used in data selection [14, 17]. This section describes procedures of estimating the corrections applied to the raw identified charged-hadron yields.
II.3.1 Reconstruction Efficiency
Figure 2: $\delta_{\rm FD}$ values as a function of $p_{T}$ for protons and antiprotons, measured in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The shaded bands indicate the systematic uncertainty.
The geometric acceptance of the detectors, detector efficiency, and analysis cuts have been taken into account by the reconstruction efficiency factor ($\epsilon_{\rm rec}$), estimated using a single-particle Monte-Carlo (MC) simulation. The MC simulation was carried out using the PHENIX integrated simulation application (PISA) package [38], which is based on GEANT3 [39] simulation software. In the PISA project the geometry, material, spatial, momentum and energy resolutions of the PHENIX detector subsystems as well as the configuration of the magnetic field are simulated in full accordance with the structure of the real facility. Simulated events have been processed using the same analysis criteria as the data. In that way, the reconstruction efficiency $\epsilon_{\rm rec}$($p_{T}$) can be estimated by the number of reconstructed hadrons ($N_{\rm rec}^{\rm MC}$) divided by the total number of hadrons generated in the MC model ($N_{\rm tot}^{\rm MC}$):
$\epsilon_{\rm rec}(p_{T})=\frac{dN_{\rm rec}^{\rm MC}/dp_{T}}{dN_{\rm tot}^{\rm MC}/dp_{T}}.$ (3)
II.3.2 Weak Decay Correction
Hyperon weak decays are an additional source of contamination for $p$ and $\bar{p}$ yields. The fraction of detected $p$ and $\bar{p}$ that come from hyperon weak decays can be taken into account by feed-down corrections $C_{\rm FD}=1-\delta_{\rm FD}$, where $\delta_{\rm FD}$ is determined as a fraction of $p$ (or $\bar{p}$) that come from hyperon weak decays to the total number of registered protons (antiprotons). The greatest contribution to the amount of decay protons and antiprotons is made by $\Lambda$ and $\bar{\Lambda}$ decays [14]. The contribution of $\Sigma^{0}$ is implicitly taken into account when considering $\Lambda$ and $\bar{\Lambda}$ decays, because $\Sigma^{0}$ decays electromagnetically with 100% branching ratio to $\Lambda$ and a photon. Other weak hyperon decays, such as $\Sigma^{\pm}$, $\Xi$, etc. have not been considered in the present analyzes. The contribution of charged $\Sigma$ states, $\Xi$ and $\Omega$ multistrange baryon states is small [17, 37] and was included in the overall systematic uncertainty estimates.
MC simulations of $\Lambda$ and $\bar{\Lambda}$ are used to calculate the fraction of $p$ and $\bar{p}$ that come from $\Lambda$ and $\bar{\Lambda}$ decays. Figure 2 shows the values of $\delta_{\rm FD}$ as a function of $p_{T}$, obtained in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions.
II.3.3 Bias-factor correction
The MB trigger is biased for events with high multiplicity, leading to miscalculations of hadron invariant yields in small collision systems [40]. The bias of the MB trigger arises mostly from single- and double-diffractive events, which result in particle production very close to the beam line. The BBC trigger is therefore biased to the nondiffractive collisions, which have larger particle production at midrapidity. In such cases, there will exist a bias towards higher charge deposition in the BBC and, hence, towards larger centrality [40]. The bias effect was corrected by applying bias-factor corrections ($f_{\rm bias}$). The values of $f_{\rm bias}$ corrections are estimated with the help of Glauber MC simulations and presented in Table 3.
II.4 Systematic Uncertainties
Systematic uncertainties are divided into three types: type A, type B, and type C. Uncertainties of type A are point-to-point uncorrelated in transverse momentum and centrality. Type B uncertainties are point-to-point correlated with transverse momentum. Type A and B uncertainties arise as a result of acceptance uncertainties, weak decay correction uncertainty, applying track selection and PID, cuts. The values of type A and B uncertainties have been estimated by varying analysis-cut parameters. Track selection and acceptance uncertainties mostly cancel for measurements of $p/\pi$ and $K/\pi$ ratios [17, 14], so only uncertainties associated with PID cuts and weak decay correction are shown with the results presented in Table 4.
Uncertainties of type C are fully correlated in transverse momentum, i.e. uncertainties of type C shift hadron yields equally in the entire range of the transverse momentum. In this work, the Type C uncertainties are attributed to the uncertainty in estimation of $\langle N_{\rm coll}\rangle$ and $f_{\rm bias}$ values.
Table 3: Summary of the $f_{\rm bias}$ correction values calculated using Glauber Monte Carlo simulation.
Collision system Centrality $f_{\rm bias}$
$p$$+$Al 0–72% 0.80$\pm$0.02
0–20% 0.81$\pm$0.01
20–40% 0.90$\pm$0.02
40–72% 1.05$\pm$0.04
${}^{3}$e$+$Au 0–88% 0.89$\pm$0.01
0–20% 0.95$\pm$0.01
20–40% 1.01$\pm$0.01
40–60% 1.02$\pm$0.01
60–88% 1.03$\pm$0.05
Table 4: Values of type A uncertainties summed quadratically to the values of type B uncertainties (%) on the identified charged hadron invariant yields and ratios. The $p_{T}$ ranges correspond to Table 2
. $p_{T}$ (GeV/$c$) System Hadron 0.5–1 1–1.5 1.5–2.0 2.0–3.0 3.0–4.0 $p$$+$Al $\pi^{+}$ 9.7 10.5 11.3 $\pi^{-}$ 8.7 11.0 10.5 $K^{+}$ 7.9 10.2 13.7 $K^{-}$ 7.7 9.9 15.0 $\bar{p}$ 9.3 7.9 7.7 8.8 $K^{+}/\pi^{+}$ 3.6 3.9 8.1 $K^{-}/\pi^{-}$ 3.6 3.7 8.7 $\bar{p}/\pi^{-}$ 6.0 6.1 6.6 6.6 ${}^{3}$e$+$Au $\pi^{+}$ 5.7 4.2 5.7 6.7 $\pi^{-}$ 11.6 11.3 10.5 9.1 $K^{+}$ 8.3 8.6 10.0 $K^{-}$ 8.7 9.9 11.3 $p$ 8.6 8.6 8.8 8.5 8.8 $\bar{p}$ 9.1 10.0 10.4 10.3 10.6 $K^{+}/\pi^{+}$ 5.1 4.0 5.1 $K^{-}/\pi^{-}$ 5.0 4.0 5.2 $p/\pi^{+}$ 7.1 6.4 7.2 7.8 $\bar{p}/\pi^{-}$ 7.2 6.4 6.5 7.9 Cu$+$Au $\pi^{+}$ 10.2 10.9 10.5 9.9 $\pi^{-}$ 8.9 10.1 10.0 10.5 $K^{+}$ 10.2 8.3 7.5 $K^{-}$ 10.0 7.7 8.6 $p$ 10.3 11.5 12.4 13.1 16.5 $\bar{p}$ 8.4 8.6 10.3 10.9 12.8 $K^{+}/\pi^{+}$ 7.0 7.0 7.7 $K^{-}/\pi^{-}$ 7.2 7.2 8.0 $p/\pi^{+}$ 8.5 8.6 10.0 10.1 $\bar{p}/\pi^{-}$ 9.1 8.5 8.7 8.7 U$+$U $\pi^{+}$ 16.8 16.5 14.7 13.7 $\pi^{-}$ 6.2 8.8 9.4 16.9 $K^{+}$ 10.2 8.3 8.3 $K^{-}$ 8.8 7.6 8.5 $p$ 10.0 9.8 9.8 11.2 14.1 $\bar{p}$ 10.3 10.0 9.1 10.5 15.2 $K^{+}/\pi^{+}$ 6.9 7.0 7.4 $K^{-}/\pi^{-}$ 8.0 7.0 7.8 $p/\pi^{+}$ 9.0 8.5 9.0 8.6 $\bar{p}/\pi^{-}$ 10.0 9.6 9.0 9.2
II.5 Invariant Spectra
The $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ invariant transverse momentum spectra were measured in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and in U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV as follows:
$\frac{1}{2\pi p_{T}}\frac{d^{2}N}{dp_{T}dy}=\frac{1}{2\pi p_{T}}\frac{N_{h}f_{\rm bias}C_{\rm FD}}{N_{\rm evt}\varepsilon_{\rm rec}\Delta p_{T}\Delta y},$ (4)
where $N_{h}$ is the raw yield of hadron $h$, $N_{\rm evt}$ is the number of nucleus-nucleus collisions, $y$ is rapidity, $C_{\rm FD}$ is the feed-down correction and $f_{\rm bias}$ is the bias-factor correction. The $C_{\rm FD}=1$ for $\pi^{\pm}$ and $K^{\pm}$; $f_{\rm bias}=1$ for the large collision systems. The resulting $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ invariant $p_{T}$ spectra are presented in Fig. 3.
The $\pi$, $K$, and $p$ invariant $p_{T}$ spectra exhibit different shapes. To quantify these differences, invariant-transverse-mass ($m_{T}=\sqrt{p_{T}^{2}+m_{0}^{2}}$) spectra were calculated. The $m_{T}$ invariant spectra of all identified charged hadrons have exponential form for $m_{T}$ $<$ 1.5 GeV and can be approximated by Eq. 5 [14]:
$\frac{1}{2\pi m_{T}}\frac{d^{2}N}{dm_{T}dy}=\frac{A}{2\pi T(T+m_{0})}\exp\left(-\frac{m_{T}-m_{0}}{T}\right)$ (5)
where $T$ is the inverse-slope parameter and $A$ is a normalization factor. Examples of resulting $\pi^{+}$, $K^{+}$ and $p$ invariant-$m_{T}$ spectra measured in central Cu$+$Au collisions are presented in Fig. 5. Approximations by Eq. 5 are shown with [red] solid lines. Values of inverse-slope parameters $T$ calculated in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions are summarized in Table 6.
Figure 5 shows examples of $T$ parameter vs. hadron mass ($m_{0}$) dependencies for different centralities of Cu$+$Au collisions. Ordering of pion, kaon, and proton inverse slope values $T_{\pi}<T_{K}<T_{p}$ can be seen in all centralities.
[…]
$T=T_{0}+m\left<u_{t}\right>^{2},$ (6)
where $T_{0}$ can be interpreted as a freeze-out temperature and $u_{t}\right>$ as the average collective velocity for all particle species.
III RESULTS
Figure 3: The $\pi^{+}$, $\pi^{-}$, $K^{+}$,$K^{-}$, $p$ and $\bar{p}$ invariant $p_{T}$ spectra measured in different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Invariant $p_{T}$ spectra are multiplied by powers of ten for clarity of presentation.
Figure 4: An example of $\pi^{+}$, $K^{+}$, $p$ invariant $m_{T}$ spectra measured in central Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV. Approximations with Eq. 5 are presented with red solid lines.
Figure 4: An example of $\pi^{+}$, $K^{+}$, $p$ invariant $m_{T}$ spectra measured in central Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV. Approximations with Eq. 5 are presented with red solid lines.
Figure 5: Mass and centrality dependence of inverse slope parameters $T$ for $\pi^{+}$, $K^{+}$ and $p$ in Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV. The dotted lines represent a linear fits of the $T(m_{0})$ values from each centrality bin by Eq. 6. The $T$ values measured for $\pi^{+}$, $K^{+}$, and $p$ in $p$$+$$p$ collisions are shown for comparison.
Figure 6: Freeze-out temperature ($T_{0}$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $T_{0}$ values measured in $p$$+$$p$ collisions are shown for comparison.
Figure 6: Freeze-out temperature ($T_{0}$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $T_{0}$ values measured in $p$$+$$p$ collisions are shown for comparison.
Figure 7: Collective velocities ($u_{t}\right>$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons in different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $u_{t}\right>$ values measured in $p$$+$$p$ collisions are shown for comparison.
Table 5: Identified charged hadron inverse slope parameters $T$ (MeV/$c^{2}$) calculated in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, U$+$U collision systems.
Table 6: Freeze-out temperatures and averaged collective velocities calculated for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions
System $\langle N_{\rm part}\rangle$ $\pi^{+}$ $\pi^{-}$ $K^{+}$ $K^{-}$ $p$ $\bar{p}$
$p$$+$Al 3.1 $\pm$ 0.1 178.88 $\pm$ 0.35 187.36 $\pm$ 0.18 210.69 $\pm$ 0.77 206.61 $\pm$ 0.78 269.51 $\pm$ 1.25
4.4 $\pm$ 0.3 183.83 $\pm$ 0.28 192.56 $\pm$ 0.29 216.10 $\pm$ 1.22 211.25 $\pm$ 1.21 275.14 $\pm$ 1.16
3.3 $\pm$ 0.1 178.20 $\pm$ 0.32 186.96 $\pm$ 0.33 210.05 $\pm$ 1.43 206.25 $\pm$ 1.46 266.62 $\pm$ 2.28
1.6 $\pm$ 0.2 173.88 $\pm$ 0.46 181.99 $\pm$ 0.26 204.74 $\pm$ 1.16 201.47 $\pm$ 1.18 254.20 $\pm$ 1.82
${}^{3}$e$+$Au 11.3 $\pm$ 0.5 208.70 $\pm$ 0.07 187.25 $\pm$ 0.06 235.84 $\pm$ 0.24 236.98 $\pm$ 0.25 295.74 $\pm$ 0.20 303.59 $\pm$ 0.24
21.1 $\pm$ 1.3 214.33 $\pm$ 0.11 191.58 $\pm$ 0.09 242.57 $\pm$ 0.38 242.91 $\pm$ 0.39 309.27 $\pm$ 0.33 317.40 $\pm$ 0.76
15.4 $\pm$ 0.9 209.84 $\pm$ 0.13 188.33 $\pm$ 0.10 237.27 $\pm$ 0.44 238.49 $\pm$ 0.46 296.44 $\pm$ 0.37 295.10 $\pm$ 0.18
9.5 $\pm$ 0.6 202.71 $\pm$ 0.15 182.66 $\pm$ 0.13 227.37 $\pm$ 0.53 229.34 $\pm$ 0.55 280.01 $\pm$ 0.44 287.12 $\pm$ 0.54
4.8 $\pm$ 0.3 191.01 $\pm$ 0.18 172.40 $\pm$ 0.15 213.44 $\pm$ 0.66 215.99 $\pm$ 0.70 254.22 $\pm$ 0.55 261.83 $\pm$ 0.66
Cu$+$Au 70.4 $\pm$ 3.0 191.72 $\pm$ 0.01 205.41 $\pm$ 0.01 249.33 $\pm$ 0.03 253.80 $\pm$ 0.03 363.65 $\pm$ 0.09 344.75 $\pm$ 0.08
154.8 $\pm$ 4.1 194.29 $\pm$ 0.12 208.57 $\pm$ 0.01 251.43 $\pm$ 0.04 255.85 $\pm$ 0.09 383.85 $\pm$ 0.13 365.18 $\pm$ 0.13
80.4 $\pm$ 3.3 192.07 $\pm$ 0.02 205.67 $\pm$ 0.02 249.14 $\pm$ 0.06 253.34 $\pm$ 0.05 357.54 $\pm$ 0.16 338.92 $\pm$ 0.15
34.9 $\pm$ 1.8 185.32 $\pm$ 0.03 197.72 $\pm$ 0.03 238.05 $\pm$ 0.09 243.35 $\pm$ 0.09 313.90 $\pm$ 0.20 306.48 $\pm$ 0.20
11.5 $\pm$ 1.8 175.24 $\pm$ 0.05 186.06 $\pm$ 0.05 222.76 $\pm$ 0.18 228.90 $\pm$ 0.18 270.86 $\pm$ 0.28 263.38 $\pm$ 0.29
U$+$U 330 $\pm$ 6 198.18 $\pm$ 0.02 205.84 $\pm$ 0.21 266.49 $\pm$ 0.08 265.80 $\pm$ 0.08 382.54 $\pm$ 0.19 374.48 $\pm$ 0.21
259 $\pm$ 7 197.28 $\pm$ 0.04 202.39 $\pm$ 0.03 269.42 $\pm$ 0.12 266.36 $\pm$ 0.11 358.24 $\pm$ 0.23 353.94 $\pm$ 0.26
65 $\pm$ 6 192.42 $\pm$ 0.06 197.74 $\pm$ 0.06 259.69 $\pm$ 0.22 257.80 $\pm$ 0.20 314.92 $\pm$ 0.32 308.57 $\pm$ 0.35
. System $\langle N_{\rm part}\rangle$ $T_{0}^{+}$ $T_{0}^{-}$ $u_{t}\right>^{+}$ $u_{t}\right>^{-}$ $p$$+$Al 3.1$\pm$0.1 167.99 $\pm$ 2.17 166.49 $\pm$ 15.29 0.28 $\pm$ 0.01 0.32 $\pm$ 0.04 4.4$\pm$0.3 171.84 $\pm$ 0.92 171.33 $\pm$ 15.96 0.29 $\pm$ 0.01 0.32 $\pm$ 0.04 3.3$\pm$0.1 168.01 $\pm$ 2.15 167.88 $\pm$ 14.30 0.28 $\pm$ 0.01 0.31 $\pm$ 0.04 1.6$\pm$0.2 164.10 $\pm$ 4.39 163.80 $\pm$ 11.20 0.27 $\pm$ 0.01 0.30 $\pm$ 0.03 ${}^{3}$e$+$Au 11.3$\pm$0.5 189.16 $\pm$ 10.22 166.27 $\pm$ 1.70 0.33 $\pm$ 0.03 0.38 $\pm$ 0.01 21.1$\pm$1.3 193.84 $\pm$ 12.42 166.44 $\pm$ 3.23 0.34 $\pm$ 0.03 0.39 $\pm$ 0.01 15.4$\pm$0.9 188.99 $\pm$ 9.82 167.59 $\pm$ 1.24 0.33 $\pm$ 0.02 0.38 $\pm$ 0.01 9.5$\pm$0.6 185.05 $\pm$ 8.57 164.60 $\pm$ 0.27 0.31 $\pm$ 0.02 0.36 $\pm$ 0.01 4.8$\pm$0.3 177.02 $\pm$ 4.99 158.33 $\pm$ 3.45 0.28 $\pm$ 0.01 0.33 $\pm$ 0.01 Cu$+$Au 70.4$\pm$3.0 154.06 $\pm$ 16.63 176.03 $\pm$ 11.95 0.46 $\pm$ 0.03 0.41 $\pm$ 0.02 154.8$\pm$4.1 149.99 $\pm$ 23.98 172.82 $\pm$ 19.75 0.48 $\pm$ 0.05 0.44 $\pm$ 0.04 80.4$\pm$3.3 157.02 $\pm$ 14.55 178.11 $\pm$ 10.20 0.45 $\pm$ 0.03 0.40 $\pm$ 0.02 34.9$\pm$2.9 161.02 $\pm$ 3.86 177.79 $\pm$ 2.35 0.40 $\pm$ 0.01 0.36 $\pm$ 0.01 11.5$\pm$1.8 160.04 $\pm$ 4.50 175.86 $\pm$ 7.56 0.34 $\pm$ 0.01 0.30 $\pm$ 0.02 U$+$U 330$\pm$6 160.03 $\pm$ 12.02 170.80 $\pm$ 13.24 0.48 $\pm$ 0.02 0.46 $\pm$ 0.03 259$\pm$7 169.04 $\pm$ 0.60 174.76 $\pm$ 2.93 0.44 $\pm$ 0.01 0.43 $\pm$ 0.01 65$\pm$6 176.03 $\pm$ 11.45 182.61 $\pm$ 9.64 0.39 $\pm$ 0.02 0.37 $\pm$ 0.02
Table 6: Freeze-out temperatures and averaged collective velocities calculated for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions
The dotted lines on Fig. 5 represent linear fits of the $T(m_{0})$ values from each centrality bin by Eq. 6. The fit parameters for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons calculated in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collision systems are shown in the Table 6 and presented in Figs. 7 and 7 as a function of $\langle N_{\rm part}\rangle$ values. Presented error values for $T_{0}^{\pm}$ and $u_{t}\right>^{\pm}$ parameters are due to the fit uncertainties. The $T_{0}$ values calculated in collisions with different geometries and centralities were found to be coincident within uncertainties, indicating that the freeze-out temperature is approximately independent of $\langle N_{\rm part}\rangle$ values. In contrast, the strength of the average transverse flow ($u_{t}\right>$ values) increases from small to large $\langle N_{\rm part}\rangle$ values, supporting the hydrodynamic picture.
Figure 8: The ratios of $p$/$\pi^{+}$ as a function of $p_{T}$ measured in different centralities of ${}^{3}$e$+$Au collisions and $\bar{p}$/$\pi^{-}$ ratios measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points, measured in $p$$+$$p$ collisions [43], are shown as a comparison.
Figure 8: The ratios of $p$/$\pi^{+}$ as a function of $p_{T}$ measured in different centralities of ${}^{3}$e$+$Au collisions and $\bar{p}$/$\pi^{-}$ ratios measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points, measured in $p$$+$$p$ collisions [43], are shown as a comparison.
Figure 9: The ratios of $p$/$\pi^{+}$ and $\bar{p}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of Cu$+$Au, and U$+$U collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
Figure 10: The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
Figure 10: The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
Figure 11: The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of Cu$+$Au, and U$+$U collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
Figure 12: Identified charged hadron nuclear-modification factors as a function of $p_{T}$ measured in central and peripheral $p$$+$Al, $d$$+$Au and ${}^{3}$e$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
Figure 12: Identified charged hadron nuclear-modification factors as a function of $p_{T}$ measured in central and peripheral $p$$+$Al, $d$$+$Au and ${}^{3}$e$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
Figure 13: Identified charged hadron nuclear-modification factors as a function of $p_{T}$ measured in central and peripheral Cu$+$Au, Au$+$Au and U$+$U collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
Figure 14: Light hadron ($\phi$ [6], $\pi^{\pm}$, $K^{\pm}$, $(p+\bar{p})/2$ and $\pi^{0}$ [44, 45]) $R_{AB}$ values vs. $p_{T}$ measured in central and peripheral Cu$+$Au, and U$+$U collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
Figure 14: Light hadron ($\phi$ [6], $\pi^{\pm}$, $K^{\pm}$, $(p+\bar{p})/2$ and $\pi^{0}$ [44, 45]) $R_{AB}$ values vs. $p_{T}$ measured in central and peripheral Cu$+$Au, and U$+$U collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
Figure 15: Comparison of $\phi$ [46, 26], $\pi^{\pm}$, $K^{\pm}$, $(p+\bar{p})/2$ and $\pi^{0}$ [47] $R_{AB}$ values vs. $p_{T}$ measured in central and peripheral $p$$+A$l and ${}^{3}$e$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
III.1 Particle Ratios
Enhancement of baryon production in nucleus-nucleus collisions is considered to be one of the signatures of QGP formation [14, 15, 17]. To investigate differences in baryon and meson production mechanisms, the ratios of $p/\pi$ and $K/\pi$ have been calculated. Figures 9 and 9 present comparisons of $p$/$\pi^{+}$ and $\bar{p}$/$\pi^{-}$ ratios in different centralities of large (Cu$+$Au, U$+$U) and small ($p$$+$$p$, $p$$+$Al, ${}^{3}$e$+$Au) collision systems. In central collisions of large systems $p$/$\pi$ ratios reach the values of $\approx$0.6, but in peripheral collisions the values of $p/\pi$ ratios are smaller than 0.4 in the whole $p_{T}$ range. Behavior of $p/\pi$ ratios observed in Cu$+$Au and U$+$U collision systems can be qualitatively described using recombination models [23, 24].
In small collision systems ($p$$+$Al, ${}^{3}$e$+$Au), the values of $p$/$\pi$ ratios are similar to those measured in $p$$+$$p$ collisions [43]. In ${}^{3}$e$+$Au collisions a modest centrality dependence can be seen, similar to that observed in $d$$+$Au collisions [14, 17]. The modest centrality dependence can be understood in terms of the small range of $\langle N_{\rm part}\rangle$ values relative the large range of $\langle N_{\rm part}\rangle$ values in large collision systems. The values of $p/\pi$ ratios measured in all centrality classes of $p$$+$Al collisions and in $p$$+$$p$ collisions are consistent within uncertainties. For all measurements type A uncertainties sum quadratically to type B uncertainties and are shown as a rectangles around the experimental points in Figs.1115.
The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ are shown in Figs. 11 and 11. The values of $K/\pi$ ratios show a modest centrality dependence, which is insignificant within systematic uncertainties. The centrality dependence of $K/\pi$ ratios in $d$$+$Au and Au$+$Au collisions was attributed to a strangeness-enhancement effect [17].
III.2 Nuclear-Modification Factors
To quantify differences of hadron production in relativistic nucleus-nucleus collisions (A+B) and in $p$$+$$p$ collisions, nuclear modification factors ($R_{AB}$) were calculated as follows:
$R_{AB}=\frac{1}{\left<N_{\rm coll}\right>}\frac{d^{2}N_{A+B}/dp_{T}dy}{d^{2}N_{p+p}/dp_{T}dy},$ (7)
where $d^{2}N_{A+B}/dp_{T}dy$ and $d^{2}N_{p+p}/dp_{T}dy$ are the invariant spectra measured in A+B and $p$$+$$p$ collisions, respectively.
Figures 13 and 13 present comparisons of identified charged-hadron $R_{AB}$ values as a function of $p_{T}$ in central and peripheral $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $R_{AB}$ values are found to be in agreement in collisions with different geometries, but with the same $\langle N_{\rm part}\rangle$ values, indicating that identified charged-hadron production depends only on system size and not geometry.
The following features of identified charged-hadron production in $p$$+$Al collisions have been found: (i) the slope of $R_{AB}(p_{T})$ in $p$$+$Al collisions is flatter than it is in ${}^{3}$e$+$Au and $d$$+$Au collisions and (ii) proton $R_{AB}$ values in $p$$+$Al collisions at the intermediate $p_{T}$ range (1.0 GeV/$c$ $<~{}p_{T}~{}<$ 2.5 GeV/$c$) are equal to unity, while in ${}^{3}$e$+$Au and $d$$+$Au collisions proton $R_{AB}$ is above unity. Differences in identified charged-hadron production between $p$$+$Al and $d/$${}^{3}$e$+$Au might be caused by the size of the $p$$+$Al system being insufficient to observe an increase in proton production.
Comparisons of identified charged-hadron $R_{AB}$ values with neutral-meson $R_{AB}$ values [47, 44, 45] in small and large collision systems are presented in Figs. 15 and 15. In large collision systems and in the ${}^{3}$e+Au collision system, proton $R_{AB}$ values are enhanced over all meson $R_{AB}$ values. The mass of the $\phi$-meson $m_{\phi}$=1019 MeV/$c^{2}$ is similar to the proton mass $m_{p}$=938 MeV/$c^{2}$, therefore the enhancement of proton $R_{AB}$ values over $\phi$-meson $R_{AB}$ values suggests differences in baryon versus meson production instead of a simple mass dependence. In $p$$+$Al collisions proton $R_{AB}$ values and $R_{AB}$ values of all measured mesons are in agreement within uncertainties, which shows zero enhancement in proton to $\phi$-meson production.
IV SUMMARY
The PHENIX experiment has measured identified charged-hadron invariant $p_{T}$ and $m_{T}$ spectra and nuclear-modification factors $R_{AB}$, and $p/\pi$, $K/\pi$ ratios in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and in U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV.
The values of freeze-out temperatures $T_{0}$ and average collective velocities $u_{t}\right>$ have been obtained. The $T_{0}$ values do not exhibit any dependence on the collision centrality and $\langle N_{\rm part}\rangle$ values, whereas the values of $u_{t}\right>$ smoothly increase with increasing of $\langle N_{\rm part}\rangle$ values. This indicates that in collisions characterized by large $\langle N_{\rm part}\rangle$ values (central Cu$+$Au, Au$+$Au, U$+$U collisions), collective effects are more pronounced than in collision systems with small $\langle N_{\rm part}\rangle$ values ($p$$+$Al, ${}^{3}$e$+$Au collisions and peripheral collisions of large systems).
The $p$/$\pi$ and $K/\pi$ ratios have been measured to investigate differences in baryon vs. meson production in large and small collision systems. In collisions with small $\langle N_{\rm part}\rangle$ values the ratios of $p$/$\pi$ are comparable to those measured in $p$$+$$p$ collisions ($(p/\pi)^{p+p}$). In collision systems with large $\langle N_{\rm part}\rangle$ values $p$/$\pi$ ratios reach the values of $\approx$0.6, which is $\approx$2 times larger than $(p/\pi)^{p+p}$.
In heavy ion collisions (Cu$+$Au and U$+$U) the $p$/$\pi$ ratios exhibit strong centrality dependence, but in small ${}^{3}$e$+$Au collisions the centrality dependence is much more modest due to the much smaller range of $\langle N_{\rm part}\rangle$ values. Within uncertainties, the $p$/$\pi$ values measured in each centrality of $p$$+$Al collisions are consistent with those measured in $p$$+$$p$ collisions. The values of the $K/\pi$ ratios show a modest centrality dependence, which is insignificant within systematic uncertainties. The observed behavior of $p/\pi$ and $K/\pi$ ratios can be qualitatively described through hadronization by recombination [23, 24].
Comparison of identified charged-hadron nuclear-modification factors shows that in $d$$+$Au, ${}^{3}$e$+$Au, Cu$+$Au, Au$+$Au, and U$+$U collision systems the $R_{AB}$ values are consistent at the same number of participant nucleons $\langle N_{\rm part}\rangle$. This indicates that identified charged-hadron production does not depend on the geometry and collision species, but rather is determined by system size alone (as indicated by $\langle N_{\rm part}\rangle$ values). Further, it is found that: (i) the slope of $R_{AB}(p_{T})$ in $p$$+$Al collisions is flatter than in ${}^{3}$e$+$Au and $d$$+$Au collisions at the same $\langle N_{\rm part}\rangle$ values, (ii) proton $R_{AB}$ values in $p$$+$Al collisions are equal to unity in the range of 1.0 GeV/$c$ $<p_{T}<$ 2.5 GeV/$c$, while proton $R_{AB}$ values measured in ${}^{3}$e$+$Au and $d$$+$Au collisions are larger than unity. Despite the observed absence of proton enhancement in $p$$+$Al collisions, there were evidences of QGP formation found by the PHENIX experiment in studies of $J/\psi$, $\psi(2S)$, and charged-hadron production at backward rapidity [48]. The observed differences in identified charged hadron production in $p$$+$Al and $d$/${}^{3}$e$+$Au collisions may be caused by the size of the system created in $p$$+$Al collisions being too small to observe the expected increase in proton production.
ACKNOWLEDGMENTS
We thank the staff of the Collider-Accelerator and Physics Departments at Brookhaven National Laboratory and the staff of the other PHENIX participating institutions for their vital contributions. We acknowledge support from the Office of Nuclear Physics in the Office of Science of the Department of Energy, the National Science Foundation, Abilene Christian University Research Council, Research Foundation of SUNY, and Dean of the College of Arts and Sciences, Vanderbilt University (U.S.A), Ministry of Education, Culture, Sports, Science, and Technology and the Japan Society for the Promotion of Science (Japan), Natural Science Foundation of China (People’s Republic of China), Croatian Science Foundation and Ministry of Science and Education (Croatia), Ministry of Education, Youth and Sports (Czech Republic), Centre National de la Recherche Scientifique, Commissariat à l’Énergie Atomique, and Institut National de Physique Nucléaire et de Physique des Particules (France), J. Bolyai Research Scholarship, EFOP, the New National Excellence Program (ÚNKP), NKFIH, and OTKA (Hungary), Department of Atomic Energy and Department of Science and Technology (India), Israel Science Foundation (Israel), Basic Science Research and SRC(CENuM) Programs through NRF funded by the Ministry of Education and the Ministry of Science and ICT (Korea). Ministry of Education and Science, Russian Academy of Sciences, Federal Agency of Atomic Energy (Russia), VR and Wallenberg Foundation (Sweden), University of Zambia, the Government of the Republic of Zambia (Zambia), the U.S. Civilian Research and Development Foundation for the Independent States of the Former Soviet Union, the Hungarian American Enterprise Scholarship Fund, the US-Hungarian Fulbright Foundation, and the US-Israel Binational Science Foundation.